OpenStudy (anonymous):

how do you solve this problem? n^3/4 =4n

OpenStudy (anonymous):

Multiply it out, get it all on one side and factor

OpenStudy (anonymous):

\[n^3\div4=4n\] try to get the n on one side

OpenStudy (anonymous):

could please show me the method? I need to see the steps to learn. Thank you.

OpenStudy (anonymous):

n^3/4 =4n -> n^3 = 16n -> n^3 - 16n = 0 OK so far?

OpenStudy (anonymous):

yes. Thank you.

OpenStudy (anonymous):

n^3 - 16n = 0 -> n(n^2-16)=0 Can you do it now?

OpenStudy (anonymous):

The next step please?

OpenStudy (anonymous):

n(n^2-16)=0 -> n(n+4)(n-4) = 0

OpenStudy (anonymous):

is it n=-4 and n= 4

OpenStudy (anonymous):

It is a cubic equation, 3 roots...

OpenStudy (anonymous):

how do I get the last root please?

OpenStudy (anonymous):

We have n(n+4)(n-4) = 0 (3 things multiplied together to get 0) Maybe if I write (n+0)(n+4)(n-4) = 0 it is clearer?

OpenStudy (anonymous):

You got it now, right?

OpenStudy (anonymous):

yes, it's n=0 n=-4 n=4 right?

OpenStudy (anonymous):

That's right.

OpenStudy (anonymous):

Thank you so much. I appreciate it.

OpenStudy (anonymous):

ur welcome.